OpenStudy (anonymous):

Find dx/dy for the equation (1/x^2)+ (1/y^2)=3 .

OpenStudy (anonymous):

x^-2 + y ^-2 = 3 -2x^-3 + -2y^-3* dy/dx = 0 2y^-3* dy/dx = -2x^-3 dy/dx = -2x^-3 / 2y^-3

OpenStudy (anonymous):

simplifying dy/dx = -y^3 / x^3

OpenStudy (anonymous):

so essentially (y)^3 - ------- (x)^3

OpenStudy (anonymous):

ohhh okay

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

hope ur curiosity 's been satisfied

OpenStudy (anonymous):

Lol. Its been more than satisfied. ;)

OpenStudy (anonymous):

okay!!!!